Jam has only 11.2g of potassium hydroxide and wants to know what mass of potassium sulfate he could produce if he reacts this with excess sulfuric acid according to the equation
2KOH(aq)+H2SO4(aq)-K2SO4(aq)+2H2O(l)

How many mole of potassium hydroxide does jam have?

How many mole of potassium sulfate can be produced?

What mass of potassium sulfate could Jam produce?
1
Expert's answer
2021-07-16T01:28:57-0400


Solution


"a.)\frac{1 mole\times11.2g KOH}{56.1056gKOH} = 0.1996 moles"


b.) Moles of K2S04

2 moles KOH : 1 mole K2S04

"\frac{0.1996 moles}{2} = 0.099 moles of K_2S0_4"


c.) "\frac{0.099moles\times174.3gK_2S0_4}{1mole} = 17.397 g"


17.397g of K2S04 will be produced